Problem 1 (Electron Beam Diffraction/De Broglie Wavelength):

A beam of 125-eV electrons is scattered from a crystal in an X-ray diffraction experiment, and a first- order peak is observed an angle of 38 from the normal to surface of the crystal. What is the spacing between planes in the diffracting crystal?

can you please show steps in solving this problem
Answer #1

energy = hc/λ

125*1.6*10^-19 = 6.626*10^-34*3*10^8 / λ

λ = 9.94*10^-9 m

using the eqation

2d sin θ = nλ

2*d* sin 38 = 1*9.94*10^-9

d = 8.07*10^-9 m

so the spacing is 8.07 nm